An object moving north with an initial velocity of 14 m/s accelerates 5 m/s2 for 20 seconds. What is the final velocity of the object?
1
Expert's answer
2020-09-28T08:10:25-0400

"v_f=v_0+at=14+5\cdot20=114(m/s)"